OpenStudy (aripotta):

t^2 - k^2 < 6 t + k > 4 If t and k are positive integers in the inequalities above and t > k, what is the value of t? a. 1 b. 2 c. 3 d. 4 e. 5

OpenStudy (anonymous):

Plug in each answer option and that's how you get the answer. @AriPotta

OpenStudy (aripotta):

i know the answer, i just don't know how they got to that particular answer and i'd like to see what other people would do. it's not for a class or anything. i'm just studying for the SAT

OpenStudy (anonymous):

t^2 - k^2 < 6 t + k > 4 based on eq1, the difference between t and k should be relevantly small (right?) when t=4 and k=3, then 4^2 - 3^2 =7 which is more than 6, and it shouldn't be lets pick, t=3 and k=2, then 3^2 - 2^2 = 5 and 5 is less than 6. Also, eq2, 3 + 2 > 4 So t=3 and k=2 When you choose smaller integers equation 2 won't work When you choose bigger integers, equation 1 won't work

OpenStudy (anonymous):

Hope I helped you with a bunch of bs that I have put up.

OpenStudy (aripotta):

lol thank you, it did help :)

OpenStudy (anonymous):

yw:)